# Tunneling time, the barrier time-delay and weak measurement What was or can actually be measured with the Attoclock?

### v1

@@ Conclusion
The tunneling time shows a universal behavior that we call UTTP, where the barrier time-delay can be defined by $\tau_{delt}\equiv\tau_{barrier}\equiv\tau_{dwell}$. It is shown that $\tau_{delt}$ agrees with the extracted time-delay $\tau_{barrier,exp}$ from the difference between adiabatic and nonadiabatic experimental measurement result of the same experiment. That is the attoclock experiment on He-atom with the adiabatic and nonadiabatic field calibration of Landsman and Hofmann Landsman et al. (2014); Hofmann et al. (2019), respectively. Our result provides conceivable definitions of the tunnel-ionization and barrier time-delays and the interpretation of the attoclock measurement. We also found that the barrier time-delay $\tau_{delt}$ corresponds to the LC-time and the interaction time Büttiker (1983); Steinberg (1995); Ramos et al. (2020); Spierings and Steinberg (2021). This is particularly evident in the limit of thick (opaque) barrier, where $\tau_{delt}$ is close to the time-delay in the nonadiabatic field calibration $\tau_{dion}$, where the back reaction is small and the weak measurement approach is justified. We assume that there is a similarity to the measurement presented in Ramos et al. (2020). In the future, our concern will be devoted to dealing with the superluminal tunnel effect, e.g. Dumont and Rivlin (2023), one of the most exciting and controversial phenomena in quantum physics.

### v2

@@ Conclusion
in the present work we showed that the tunneling time shows universal behavior UTTP, where the barrier time delay can be convincingly defined by $\tau_{delt}\equiv\tau_{barrier}\equiv\tau_{dwell}$. It is shown that $\tau_{delt}$ agrees with time delay $\tau_{barrier,exp}$, the difference between adiabatic and nonadiabatic field calibrations of the measurement resuls of the attoclock experiment on He-atom of Landsman and Hofmann Landsman et al. (2014); Hofmann et al. (2019), respectively. The result provides conceivable definitions of the tunnel ionization and barrier time delays and the interpretation of the attoclock measurement. We also found that the barrier time delay $\tau_{delt}$ corresponds to the LC-time and the interaction time Büttiker (1983); Steinberg (1995); Ramos et al. (2020); Spierings and Steinberg (2021). This is particularly evident in the limit of thick barrier, where $\tau_{delt}$ is closer to the time delay in the nonadiabatic field calibration $\tau_{dion}$, and where the back reaction is small and the weak measurement approach is justified. We assume that there is a similarity to the measurement presented in Ramos et al. (2020). In the future, our concern is devoted to the superluminal tunneling, e.g. Dumont and Rivlin (2023), one of the most exciting phenomena in quantum physics.